OpenStudy (anonymous):

find the indefinite integral ∫ 1/x^2 (e ^-1/x) dx

OpenStudy (anonymous):

The ambiguity kills... \[\huge \int\limits_{}^{}\frac{1}{x^2(e^{-\frac{1}{x}})}dx\] Is this it?

OpenStudy (anonymous):

nope the (e ^-1/x) is not in the denominator. but it's right :)

OpenStudy (anonymous):

It's not in the denominator, but it's right? It's wrong!!!! But seriously.. :D \[\huge \int\limits_{}^{}\frac{e^{-\frac{1}{x}}}{x^2}dx\] This one then?

OpenStudy (anonymous):

i guess it could be placed that way.. but then it's kinda just multiplied to it :))

OpenStudy (anonymous):

Okay, since you're so particular :P \[\huge \int\limits_{}^{}\frac{1}{x^2}(e^{-\frac{1}{x}})dx \] This one better? :D

OpenStudy (anonymous):

perfect. haha now how do we answer that ? haha

OpenStudy (anonymous):

Well, always look for a way to simplify things... Too bad there is no clear-cut way to integrate, it's not like trying to find the derivative, where you always know what you're doing, this takes a bit of guesswork :D So... is there any bit of that integrand (fancy name for the function being integrated) whose derivative is also in the integrand?

OpenStudy (anonymous):

My rule of thumb here is... always suspect the most complicated looking function :D

OpenStudy (anonymous):

Stumped? Let me know :)

OpenStudy (anonymous):

Suspicious looking complicated function. That's actually a pretty good advice! (-: \[\Large \checkmark \]

OpenStudy (anonymous):

:D I just thought it up now... :/

OpenStudy (anonymous):

So it has only been a rule of thumb for a few minutes, thanks though :D @alyannahere Are you still here?

OpenStudy (anonymous):

here ! sorry. i think the e ^-1/x looks pretty complicated

OpenStudy (anonymous):

Cool, I thought so :P

OpenStudy (anonymous):

i think i get it ? but then i'm not sure how it works.. like i know there s a property but i can't seem to apply it

OpenStudy (anonymous):

Go, you can tell me anything :D I don't guarantee I'll liste..... LOL What was your solution? :)

OpenStudy (anonymous):

there's this property stating... ∫e^f(x) X f'(x) X dx = e^f(x) + C X is multiplied

OpenStudy (anonymous):

Let me try to rid myself of ambiguities again.... \[\huge \int\limits e^{f(x)}Xf'(x)Xdx = e^{f(x)}+C\] I don't quite get it... what's with the big X's? :/

OpenStudy (anonymous):

X is multiply :)) haha times :))

OpenStudy (anonymous):

/head scratching/ \[\huge \int\limits e^{f(x)}f'(x)dx = e^{f(x)} + C\] is it?

OpenStudy (anonymous):

yup! that's one of the general indefinite integral formulas

OpenStudy (anonymous):

You know how to apply it?

OpenStudy (anonymous):

Maybe you can simply let u = -1/x du = 1/(x^2) dx Then, the intergral becomes pretty easy

OpenStudy (anonymous):

@RolyPoly how did du become 1/(x^2) dx?

OpenStudy (anonymous):

Simply take the derivative of -1/x

OpenStudy (anonymous):

\[\huge \frac{d}{dx}x^k=kx^{k-1}\],

OpenStudy (anonymous):

And you know, \[\huge -\frac{1}{x}=-x^{-1}\]is no exception :P

OpenStudy (anonymous):

for k \(\ne\) 0, I guess?

OpenStudy (anonymous):

Always missing those things :/ Yeah, for a nonzero k.

OpenStudy (anonymous):

i still don't get it

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@alyannahere Do you know how to differentiate 1/x ?

OpenStudy (anonymous):

Use this formula instead, to be more general: \[ \huge \frac{d}{dx}ax^k=akx^{k-1}\]

OpenStudy (anonymous):

oooh. i see i get it now... thank you guysss!!!!!!!!!!!

OpenStudy (anonymous):

Try to solve the problem. If you have any questions, please let us know :)

OpenStudy (anonymous):

Well, I'm bored, I'm going to play a bit with this... \[\huge \int\limits\limits e^{f(x)}f'(x)dx\]Let \[\large u=f(x)\]\[\large du = f'(x) dx\] \[\huge \int\limits\limits e^{f(x)}f'(x)dx = \int\limits e^udu\]\[\huge \int\limits e^u du = e^u + C\]Resusbstituting... \[\huge e^{f(x)} + C\] Voila D

OpenStudy (anonymous):

i did. the answer is e^-1/x +C ;)

OpenStudy (anonymous):

right? haha

OpenStudy (anonymous):

Yup! Congrats :)

OpenStudy (anonymous):

Nice work :)

OpenStudy (anonymous):

how about this problem @RolyPoly find the antiderivative of dp/dx = e^x + e^-x / (e^x - e^-x)^2 ?

OpenStudy (anonymous):

well i made u = e^x + e^-x du = e^x - e^-x then i got stuck

OpenStudy (anonymous):

Basically trying to integrate... \[\huge \int\limits \frac{e^x + e^{-x}}{(e^x-e^{-x})^2}\] OKAAAY Second rule that I just made up... If you got two functions that are both suspiciously complicated, go for the one in the denominator... that's probably it :P

OpenStudy (anonymous):

@alyannahere That's correct. Then, what will the integral become?

OpenStudy (anonymous):

@RolyPoly Are you sure it's correct? I'd let u = e^x - e^-x

OpenStudy (anonymous):

hahahaah i like your rules @PeterPan & yup! :)

OpenStudy (anonymous):

well @RolyPoly it will become u (u^-2)

OpenStudy (anonymous):

@PeterPan 100% sure. The asker let u = e^x + e^(-x), not u = e^x - e^(-x)